We know that Costco’s free cash flow (FCF) is $3,330 million, the financial liability is $22,977 million, the liquid asset is $7,259 million, the shares outstanding is 440 million, and the stock price per share is $302.62. According to the consensus of financial analysts following Costco, the WACC is 8%, and the 5-year growth rate of FCF is about 9.3%. What is the market’s expectation for the long-term growth rate of Costco’s FCF?

A. 5.17% B. 4.80% C. 6.34% D. 5.05%
Answer #1

hence, the correct answer is option C) 6.34%

Costco's enterprise value = Equity value + Financial liability - liquid asset = N x P + 22,977 - 7,259 = 440 x 302.62 + 22,977 - 7,259 = $  148,870.80 million

Also enterprise value = PV of all the future FCFF = C1 / (1 + r) + C2 / (1 + r)2 + C3 / (1 + r)3 + C4 / (1 + r)4 + + C5 / (1 + r)5 + HV / (1 + r)5 = PV of FCFF over first five years + PV of horizon value

Please see the table below:

Year FCFF growth PV factor PV of FCFF
n A g (1 + 8%)^(-n) A x PV factor
0         3,300.00               1.0000
1         3,606.90 9.30%               0.9259           3,339.72
2         3,942.34 9.30%               0.8573           3,379.92
3         4,308.98 9.30%               0.7938           3,420.61
4         4,709.71 9.30%               0.7350           3,461.78
5         5,147.72 9.30%               0.6806           3,503.45
Total        17,105.48

Hence, enterprise value = 148,870.80 = PV of FCFF over first five years + PV of horizon value =  17,105.48 + PV of horizon value

Hence, PV of HV5 = 148,870.80 -  17,105.48 =  131,765.32

HV5 x (1 + 8%)-5 =  131,765.32

Hence, HV5 =  131,765.32 x 1.085 = $ 193,606.48

Hence, C5 x (1 + gt) / (r - gt) =   193,606.48

Or, 3,503.45 x (1 + gt) / (0.08 - gt) = 193,606.48

Hence, (1 + gt) / (0.08 - gt) = 55.2617

Hence, gt = (0.08 x 55.2617 - 1) / (1 + 55.2617) = 6.3%

hence, the correct answer is option C) 6.34%